# Benchmarking LLM Privacy Recognition for Social Robot Decision Making

### v2

@@ 1. Introduction
As robots become more competent and provide greater utility, we increasingly rely on them to perform tasks in private locations such as homes (New York State Office for the Aging, 2025) and care settings  (Jack, 2024; The Sun Tech, 2025). As early as 2014, more than 3,000 Paro therapeutic robots had been sold across 30 countries (of Advanced Industrial Science and, AIST), and by 2024, over 50 million Roomba units had been sold globally (iRobot Corporation, 2025). In 2024, the market size for consumer robots increased to \$10.92 billion, with a forecast of \$40.15 billion in 2030 (Grand View Research, 2024).
@@
Like many other technologies, robots pose a threat to users’ data privacy (Guo, 2022). Robots can autonomously navigate human environments and collect data in private spaces that users may not have originally intended or desired. Beyond traditional data privacy concerns that robots pose as a connected device collecting audio or visual data (Fell, 2024), they introduce a new threat to users’ interpersonal privacy (Tang et al., 2022; Bell et al., 2025). Robots can interact with users in a human-like manner, encouraging engagement and communication. Once a robot has captured user data in this interpersonal context, it can then potentially share that data with other users. This phenomenon is unique to robots in home settings, placing them in a critical position to engage in privacy decision making. It is within this unique context that large language models (LLMs) may have the potential to alleviate privacy concerns in home settings. LLMs may be able to determine whether a given piece of information is private, how that information should be processed, and with whom it may be shared. As privacy can be highly contextual (Nissenbaum, 2004, 2009), LLMs might help robots navigate the nuances of any given scenario.
@@
Simultaneously, researchers in academia and industry are deploying LLMs and video language models (VLMs) on robots. These models equip robots with the tools to perceive and reason about their environments far beyond their previous capabilities. Researchers have demonstrated that LLMs can enhance robot capabilities by enabling natural language requests (Ahn et al., 2022), writing code to control robot movement on-the-fly (Liang et al., 2023), achieving interactive dialogue (Shinn et al., 2023), and improving navigation abilities (Hassan et al., 2024). Several companies have also introduced projects that utilize LLMs or VLMs to enable more intelligent robots, such as Google’s DeepMind-RT-2 (Brohan et al., 2023) and NVIDIA’s GR00T N1 (Bjorck et al., 2025). Additionally, researchers have evaluated LLM-powered robots through user studies to identify the potential benefits and weaknesses of these LLM frameworks (Yuan et al., 2024a; Murali et al., 2023; Kim et al., 2024a; Azeem et al., 2024). From these advancements, it is clear that LLMs and VLMs can offer significant benefits when integrated with robots.
@@
As LLMs are increasingly integrated into home robots, their ability to navigate nuanced privacy decisions regarding data collection, processing, and sharing becomes critical. Prior works have primarily focused on general LLM safety, such as LLMs’ vulnerability against jailbreak attacks (Zhuo et al., 2023). Few works, however, have investigated the privacy perceptions of LLMs and VLMs (Mireshghallah et al., 2023; Shao et al., 2024), and none, to our knowledge, have done so in the context of human-robot interaction (HRI). Given this gap, we investigate how LLMs help robots interpret privacy and make privacy decisions regarding data collection, processing, and sharing in home settings. Specifically, we raise the following research questions: RQ1: How does an individual’s privacy orientation influence their privacy expectations of robots in home settings?; RQ2: How well do state-of-the-art LLMs align with individuals’ privacy expectations?; and RQ3: How can we improve state-of-the-art LLMs’ conformity with individuals’ privacy expectations?
@@
  • We discuss the design implications of LLM-powered robots, suggesting the potential of using LLMs as privacy controllers for robots. By promoting data collection indicators and adopting users’ privacy preferences collected through interaction, robots can enable context- and privacy-aware behavior tailored to an individual’s needs.
@@ 2.1. Conceptual Approaches to Privacy
This work utilizes Contextual Integrity (CI) (Nissenbaum, 2004, 2009) as its primary theoretical underpinning. This framework proposes that privacy is composed of dimensions including contexts, actors, attributes, transmission principles, and purposes. These dimensions are critical to understanding the contextual dependence of what makes a given situation private and how individuals perceive technology to be a threat to their privacy. Existing works have used this framework for tasks such as privacy-policy analysis (Sadeh et al., 2013; Amos et al., 2021) and identification of privacy norms in smart homes (Apthorpe et al., 2018). Brause and Blank (2024) have additionally proposed expanding this model to include “information consequences”, privacy skills and awareness, and choice. In this work, we consider only the original dimensions of CI in our analysis of LLM privacy awareness.
Beyond this broad framework of privacy, many other works have attempted to assess the specific privacy attitudes, orientations, and personas of individuals as they engage with technology (Preibusch, 2013). These efforts have focused on online privacy (Elueze and Quan-Haase, 2018; Buchanan et al., 2007), out-of-device privacy (i.e., privacy violations through direct human observation) (Farzand et al., 2024), privacy perceptions broadly (Chignell et al., 2003; Giang et al., 2023), and the identification of privacy personas (Hrynenko and Cavallaro, 2024). Of particular note to this work is the Privacy Orientation Scale (POS) (Baruh and Cemalcılar, 2014), which captures an individual’s perceptions of privacy as a right, concerns about their own informational privacy, other-contingent privacy, and concerns about the privacy of others. These four subscales can then be used to classify a user’s privacy orientation as fitting one of three profiles, including privacy advocates, privacy individualists, and privacy indifferents. Other works have described the general challenge of defining privacy personas (Biselli et al., 2022). Given the substantial contextual dependence of privacy, domain-agnostic assessments of privacy perceptions may also struggle to generate consistent personas across domains.
@@ 2.2. Privacy in Robotics
While privacy has been explored in a myriad of fields, it is a growing area of concern in robotics (Denning et al., 2009; Guerrero, 2017; Rueben et al., 2018; Torresen, 2018; Pagallo et al., 2018; Lutz et al., 2019; Chatzimichali et al., 2020). Particular emphasis has been placed on robotics usage in homes (Levinson et al., 2024b, a; Fernandes et al., 2016; Urquhart et al., 2019), as these environments are highly private, contain substantial amounts of sensitive information, and offer inhabitants a sense of security. Given the capacity of some robots to capture data, autonomously navigate environments, and socially engage with humans, they pose a significant threat to privacy with unique challenges compared to many other technologies. Existing works have also focused on the inherent privacy-utility trade-off that exists when robots are introduced into new environments (Butler et al., 2015; Leite and Lehman, 2016). While users may have concerns about robot privacy violations, these may be tolerated if sufficient utility is provided by the robot (Lutz and Tamó-Larrieux, 2020).
Though these concerns are widespread, many prior works have proposed solutions that attempt to limit data collection, retention, and exposure (Sullivan and Mutlu, 2025). For example, low-resolution sensors (Kim et al., 2019) and purposeful sensor selection (Eick and Antón, 2020) have been proposed to limit unnecessary data collection, while sensor redirection (Shome et al., 2023) and privacy controllers (Tang et al., 2022) have been proposed to mitigate data exposure. While these efforts have uncovered new methods for preserving human privacy, the combination of LLMs and robots presents novel opportunities for privacy-preserving LLM decision making.
@@ 2.3. Robots & LLM Integration
While mainstream LLMs are a relatively recent development (Vaswani et al., 2017; Radford et al., 2018), LLM integration in robots has been a clear area of interest for HRI applications. Though this intersection of technologies is still in its infancy, several works have already utilized LLMs in robotics applications. These applications have been specialized, such as a chemistry assistant robot (Yoshikawa et al., 2023) and a health attendant robot (Kim et al., 2024b), as well as broad, such as a ‘robotic brain’ to manage memory and control (Mai et al., 2023). Other works have considered the potential of LLM-powered robots compared to other forms of LLMs (i.e., text- and voice-based) (Kim et al., 2024a). All of these works, however, only begin to explore the full potential of robot-LLM integration. Further study is needed to understand the implications of LLMs for perception, control, interaction, and, most relevant to this work, decision making (Zeng et al., 2023).
@@ 2.4. LLM Privacy Risks & Benchmarking
Prior works have studied various LLM privacy risks and have shown that the models can leak sensitive data by extracting their memory. Specifically,  Carlini et al. (2021) have shown that a practical extraction attack could be used to recover GPT-2 training sequences. Additional works have demonstrated that LLMs can infer sensitive personal information (Kim et al., 2023; Staab et al., 2023) and remain vulnerable to Membership Inference Attacks (Duan et al., 2024; Mattern et al., 2023; Fu et al., 2023).
To assess the potential risks and benefits that LLMs present, researchers utilize benchmarking as a standardized evaluation method for LLM performance. Common evaluation domains include reasoning (Srivastava et al., 2022; Hendrycks et al., 2020; Yi et al., 2025), multilingualism (Hu et al., 2020; Liang et al., 2020), coding (Chen et al., 2021; Austin et al., 2021), and instruction following (Zheng et al., 2023; Dubois et al., 2025) as well as safety and privacy (Ruan et al., 2023; Yuan et al., 2024b; Zhang et al., 2023; Gu et al., 2024; Zhang et al., 2024b; Zharmagambetov et al., 2025; Zhu et al., 2024; Zhang et al., 2024a; Chang et al., 2025; Bagdasarian et al., 2024). PURE (Chi et al., 2023), PolicyQA (Ahmad et al., 2020), PrivacyQA (Ravichander et al., 2019), and GenAIPABench (Hamid et al., 2023) have been performed to evaluate privacy-policy comprehension using dedicated datasets. Of particular relevance to this work, Mireshghallah et al. (2023) first raised the question of whether LLMs can reason about CI. These authors utilized their benchmark, ConfAlde, to show that GPT-4 violates CI norms in 39% of evaluated scenarios where humans would withhold information. Most recently, PrivacyLens (Shao et al., 2024) introduced vignette-based questions to survey LLMs’ privacy-norm awareness. However, existing works study whether an LLM can decide on the appropriateness of sharing data using tools from CI. In this work, we aim to explore how these models make decisions, and how well their responses align with human preferences in household HRI contexts.
@@ 3.1. Dimensions of Contextual Integrity
We first reviewed existing literature on general in-home privacy concerns (Choe et al., 2011) and Contextual Integrity (Nissenbaum, 2004). From this literature, we became oriented with a comprehensive list of the many in-home activities individuals wish to keep private. Additionally, we distilled the primary dimensions of contextual privacy, including context, actors, attributes, transmission principles, and purposes. From each of these dimensions, we further defined subdimensions relevant to privacy in HRI. These expanded dimensions include type of location, robot involvement, number of users, age of users, relation of users, medium of data, protected status of data, information flow, and task dependence. Table 1 presents a full list of all original dimensions, expanded dimensions, and factors.
While most of these dimensions and their factors are fairly intuitive, the factors of protected status and medium of data are worth discussing further. Our original intent behind protected status was to capture the sensitivity of the information involved in a given scenario. Given how subjective sensitivity may be, we chose to instead define protected status as either protected or non-protected based upon legal norms. For example, information we consider to be protected includes health data (for Civil Rights, 2003), academic records (U.S. Congress, 1974), and legal records (United States Courts, 2008), among others (Internal Revenue Service, 1976; U.S. Congress, 1999). Conversely, we consider information such as vacation plans or intimacy between partners to be non-protected. Additionally, we considered only specific mediums of data that we believed an LLM or VLM could reliably recognize. These data types include visuals of the human body in compromising situations (e.g., bathing, toileting, or intimacy), visuals of device screens, audio of conversations, and visuals of physical documents. These broad categories were selected due to the wide variety of information each of them can convey.
@@ 3.2. Identification of Household Patterns
With our scenario dimensions defined, we next sought to understand the frequency with which certain scenario elements would occur in a real home setting. For example, we questioned how frequently users are at home alone or with others present (i.e., number of users). Given the contextual dependence of some dimensions on a robot’s presence, we evaluated only the numbers of users, types of locations, ages of users, and relations of users. To achieve this goal, we conducted a large online user study.
@@ 3.2.2. Participants
To conduct this study, we hosted our Qualtrics survey on Prolific. Through this platform, we recruited 300 participants who were all located within the U.S. and fluent in English. Participants ranged in age from 18 to 83 years $(M=45.88,SD=15.95)$, and the population was split 51.33% women, 48.00% men, and 0.67% non-binary. Participants received \$0.60 USD for completing the four-minute survey. To ensure that our data was representative of households within the U.S., we additionally screened participants through Prolific to recruit a sample population consistent with the U.S. census in terms of ethnicity, sex, and age.
@@ 3.2.3. Results
The results of this study can be found in Appendix D. These values are used as weights within our scenario development to determine the frequency with which specific household characteristics are represented. For all dimensions that we did not evaluate, we represent factors equally.
@@ 3.3. Scenario Generation
After defining our scenario dimensions and crowdsourcing frequency weights, we next generated our privacy scenarios. Similar to the scenario scripts developed by Tang et al. (2022), we formatted ours to include location, characters, contextual information to describe the scene, and, in most cases, a scripted dialogue between characters. Using the dimensions defined above, we first created a set of all possible combinations of dimensions and then sampled 50 combinations utilizing our crowdsourced weights. Given our sample of dimension combinations, we prompted various ChatGPT 4o, Gemini 2.0Pro, and Claude Sonnet 3.7 to generate base, privacy-relevant scenarios. We used these base scenarios to construct a scenario that properly represents the dimensions in each dimension combination. We also ensured scenarios do not have explicit references to privacy, such that neither the participants nor the studied LLMs would be primed. We present an example script in Figure 1, and the full dataset can be accessed in the OSF repository [^1] https://osf.io/wbp37/?view_only=4a2799ee7cf04480a54d06faf342714d.

### v3

@@ 1. Introduction
As robots become more competent and provide greater utility, we increasingly rely on them to perform tasks in private locations such as homes (nysofa_elliq_initiative_2025) and care settings  (Jack2024_PARIS; TheSun2025_NuraBot). As early as 2014, more than 3,000 Paro therapeutic robots had been sold across 30 countries (AIST_2013_Paro), and by 2024, over 50 million Roomba units had been sold globally (irobot2024financial). In 2024, the market size for consumer robots increased to \$10.92 billion, with a forecast of \$40.15 billion by 2030 (grandview2025consumer).
@@
Like many other technologies, however, robots pose a threat to users’ data privacy (Guo2022RoombaPrivacy). Robots can autonomously navigate human environments and collect data in private spaces that users may not have originally intended or desired. Beyond traditional data privacy concerns that robots pose as a connected device collecting audio or visual data (fell2024insecure), they introduce a new threat to users’ interpersonal privacy (tang2022confidant; bell2025always). Robots can interact with users in a human-like manner, encouraging engagement and communication. Once a robot has captured user data in this interpersonal context, it can then potentially share that data with other users. This phenomenon is particularly noteworthy in home settings where robots are placed in critical positions to engage in privacy decision making. It is within this unique context that large language models (LLMs) may hold the potential to alleviate privacy concerns. LLMs may be able to determine whether information is private, how that information should be processed, and with whom it may be shared. As privacy is highly contextual (nissenbaum2004privacy; nissenbaum2009privacy), LLMs might help robots navigate the nuances of any given scenario.
@@
While efforts are nascent, researchers in academia and industry are deploying LLMs and video language models (VLMs) on robots. These models equip robots with the tools to perceive and reason about their environments far beyond their previous capabilities. Researchers have demonstrated that LLMs can enhance robot capabilities by enabling natural language requests (ahn2022can), writing code to control robot movement on-the-fly (liang2023code), achieving interactive dialogue (shinn2023reflexion), and improving navigation abilities (hassan2024integrating). Several companies have also introduced projects that utilize LLMs or VLMs to enable more intelligent robots, such as Google’s DeepMind-RT-2 (brohan2023rt) and NVIDIA’s GR00T N1 (bjorck2025gr00t). Additionally, researchers have evaluated LLM-powered robots through user studies to identify the potential benefits and weaknesses of these LLM frameworks (yuan2024towards; murali2023improving; kim2024understanding; azeem2024llm). From these advancements, it is clear that LLMs and VLMs can offer significant benefits when integrated with robots.
@@
Though many state-of-the-art LLMs have been designed with some level of privacy protection in mind, it is not yet clear the degree to which these models are privacy aware. Prior works have primarily focused on general LLM safety, such as LLMs’ vulnerability against jailbreak attacks (zhuo2023red). Few works, however, have investigated the privacy perceptions of LLMs and VLMs (mireshghallah2023can; shao2024privacylens), and none, to our knowledge, have done so in the context of human-robot interaction (HRI). Given this gap, we investigate how LLMs may help robots interpret privacy and make privacy decisions regarding data collection, processing, and sharing in home settings. Specifically, we raise the following research questions: RQ1: How does an individual’s privacy orientation influence their privacy expectations of robots in home settings?; RQ2: How well do state-of-the-art LLMs align with individuals’ privacy expectations?; and RQ3: How can we improve state-of-the-art LLMs’ conformity with individuals’ privacy expectations?
@@
  • We discuss design implications for LLM-powered robots, suggesting the potential of using LLMs as privacy controllers. By promoting data collection indicators and adopting users’ privacy preferences collected through interaction, robots can enable context- and privacy-aware behavior tailored to an individual’s needs.
@@ 2.1. Conceptual Approaches to Privacy
This work utilizes Contextual Integrity (CI) (nissenbaum2004privacy; nissenbaum2009privacy) as its primary theoretical underpinning. This theory proposes that privacy is composed of dimensions including contexts, actors, attributes, transmission principles, and purposes. These dimensions are critical for understanding the contextual dependence of what makes a situation private and how individuals perceive technology to be a threat to their privacy. Existing works have used this theory for tasks such as privacy-policy analysis (sadeh2013usable; Amos_2021) and identification of privacy norms in smart homes (apthorpe2018discovering). Brause and Blank (brause2024there) have additionally proposed expanding this model to include “information consequences”, privacy skills and awareness, and choice. In this work, we consider only the original dimensions of CI in our analysis of LLM privacy awareness.
Beyond this broad theory of privacy, many other works have attempted to assess the specific privacy attitudes, orientations, and personas of individuals as they engage with technology (preibusch2013guide). These efforts have focused on online privacy (elueze2018privacy; buchanan2007development), out-of-device privacy (i.e., privacy violations through direct human observation) (farzand2024out), privacy perceptions broadly (chignell2003privacy; giang2023privacy), and the identification of privacy personas (hrynenko2024identifying). Of particular note to this work is the Privacy Orientation Scale (POS) (baruh2014more), which captures an individual’s perceptions of privacy as a right, concerns about their own informational privacy, other-contingent privacy, and concerns about the privacy of others. These four subscales can then be used to classify a user’s privacy orientation as fitting one of three profiles, including privacy advocates, privacy individualists, and privacy indifferents. Other works have described the general challenge of defining privacy personas (biselli2022challenges). Given the substantial contextual dependence of privacy, domain-agnostic assessments of privacy perceptions may also struggle to generate consistent personas across domains.
@@ 2.2. Privacy in Robotics
While privacy has been explored in a myriad of fields, it is a growing area of concern in robotics (denning2009spotlight; guerrero2017cybersecurity; rueben2018themes; torresen2018review; pagallo2018rise; lutz2019privacy; chatzimichali2020toward). Particular emphasis has been placed on robotics usage in homes (levinson2024snitches; levinson2024our; fernandes2016detection; urquhart2019responsible), as these environments are highly private, contain substantial amounts of sensitive information, and offer inhabitants a sense of security. Given the capacity of some robots to capture data, autonomously navigate environments, and socially engage with humans, they pose a significant threat to privacy with unique challenges compared to many other technologies. Existing works have also focused on the inherent privacy-utility trade-off that exists when robots are introduced into new environments (butler2015privacy; leite2016robot). While users may have concerns about robot privacy violations, these may be tolerated if sufficient utility is provided by the robot (lutz2020robot).
Though these concerns are widespread, many prior works have proposed solutions that attempt to limit data collection, retention, and exposure (sullivan2025protecting). For example, low-resolution sensors (kim2019privacy) and purposeful sensor selection (eick2020enhancing) have been proposed to limit unnecessary data collection, while sensor redirection (shome2023robots) and privacy controllers (tang2022confidant) have been proposed to mitigate data exposure. While these efforts have uncovered new methods for preserving human privacy, the combination of LLMs and robots presents novel opportunities for privacy-preserving LLM decision making.
@@ 2.3. Robots & LLM Integration
While mainstream LLMs are a relatively recent development (vaswani2017attention; radford2018improving), LLM integration in robots has been a clear area of interest for HRI applications. Though this intersection of technologies is still in its infancy, several works have already utilized LLMs in robotics applications. These applications have been specialized, such as a chemistry assistant robot (yoshikawa2023large) and a health attendant robot (kim2024framework), as well as broad, such as a ‘robotic brain’ to manage memory and control (mai2023llm). Other works have considered the potential of LLM-powered robots compared to other forms of LLMs (i.e., text- and voice-based) (kim2024understanding). All of these works, however, only begin to explore the full potential of robot-LLM integration. Further study is needed to understand the implications of LLMs for perception, control, interaction, and, most relevant to this work, decision making (zeng2023large).
@@ 2.4. LLM Privacy Risks & Benchmarking
Prior works have studied LLM privacy risks and shown that the models can leak sensitive data by extracting their memory. Specifically, Carlini et al. (carlini2021extracting) have shown that a practical extraction attack could be used to recover GPT-2 training sequences. Additional works have demonstrated that LLMs can infer sensitive personal information (kim2023propile; staab2023beyond), and remain vulnerable to prompt injection (costa2025securing; greshake2023not) and Membership Inference Attacks (duan2024membership; mattern2023membership; fu2023practical).
To assess the potential risks and benefits that LLMs present, researchers utilize benchmarking as a standardized evaluation method for LLM performance. Common evaluation domains include reasoning (srivastava2022beyond; yi2025privacy), multilingualism (hu2020xtreme; liang2020xglue), coding (chen2021evaluating; austin2021program), and instruction following (zheng2023judging; dubois2025lengthcontrolledalpacaevalsimpleway) as well as safety and privacy (ruan2023identifying; zhang2023safetybench; gu2024mllmguard; zhang2024multitrust; zharmagambetov2025agent; zhu2024privauditor; zhang2024multi; chang2025keep; bagdasarian2024airgapagent). PURE (chi2023plue), PolicyQA (ahmad2020policyqa), PrivacyQA (ravichander2019question), and GenAIPABench (hamid2023genaipabench) have been performed to evaluate privacy-policy comprehension using dedicated datasets. Mireshghallah et al. (mireshghallah2023can) first raised the question of whether LLMs can reason about CI. These authors utilized their benchmark, ConfAlde, to show that GPT-4 violates CI norms in 39% of evaluated scenarios where humans would withhold information. Most recently, PrivacyLens (shao2024privacylens) introduced vignette-based questions to survey LLMs’ privacy-norm awareness. However, existing works study whether an LLM can decide on the appropriateness of sharing data using tools from CI. In this work, we aim to explore how these models make decisions, and how well their responses align with human preferences in household HRI contexts.
@@ 3.1. Dimensions of Contextual Integrity
We first reviewed existing literature on general in-home privacy concerns (choe2011living) and Contextual Integrity (nissenbaum2004privacy). From this literature, we became oriented with a comprehensive list of the many in-home activities individuals wish to keep private. Additionally, we distilled the primary dimensions of contextual privacy, including context, actors, attributes, transmission principles, and purposes. From each of these dimensions, we further defined subdimensions relevant to privacy in HRI (e.g., elements that a robot may be able to identify or reason about, or that relate to its role). These expanded dimensions include type of location, robot involvement, number of users, age of users, relation of users, medium of data, protected status of data, information flow, and task dependence. Table 1 presents a full list of all original dimensions, expanded dimensions, and factors.
While most of these dimensions and their factors are fairly intuitive, the factors of protected status and medium of data are worth discussing further. Our original intent behind protected status was to capture the sensitivity of the information involved in a given scenario. Given how subjective sensitivity may be, we chose to instead define protected status as either protected or non-protected based on legal norms. For example, information we consider to be protected includes health data (office2003summary), academic records (ferpa), and legal records (fre502), among others (irc6103; glba). Conversely, we consider information such as vacation plans or intimacy between partners to be non-protected. Additionally, we considered only specific mediums of data that we believed an LLM or VLM could reliably recognize. These data types include visuals of the human body in compromising situations (e.g., bathing, toileting, or intimacy), visuals of device screens, audio of conversations, and visuals of physical documents. These broad categories were selected due to the wide variety of information each of them can convey.
@@ 3.2. Identification of Household Patterns
With our scenario dimensions defined, we next sought to understand the frequency with which certain scenario elements would occur in a real home setting. For example, we questioned how frequently users are at home alone or with others present (i.e., number of users). Given the contextual dependence of some dimensions on a robot’s presence, we evaluated only the numbers of users, types of locations, ages of users, and relations of users. To achieve this goal, we conducted an online user study with 300 participants.
@@ 3.2.2. Participants
To conduct this study, we hosted our Qualtrics survey on Prolific. Through this platform, we recruited 300 participants who were all located within the U.S. and fluent in English. Participants ranged in age from 18 to 83 years $(M=45.88,SD=15.95)$, and the population was split 51.33% women, 48.00% men, and 0.67% non-binary. Participants received \$0.60 USD for completing the four-minute survey. To ensure that our data was representative of U.S. households, we additionally screened participants through Prolific to recruit a sample population consistent with the U.S. census in terms of ethnicity, sex, and age.
@@ 3.2.3. Results
The results of this study can be found in Appendix C. These values are used as weights within our scenario development do properly represent household characteristics. For all dimensions that we did not evaluate, we represent factors equally.
@@ 3.3. Scenario Generation
After defining our scenario dimensions and crowdsourcing frequency weights, we next generated our privacy scenarios. Similar to the scenario scripts developed by Tang et al. (tang2022confidant), we formatted ours to include location, characters, contextual information to describe the scene, and, in most cases, a scripted dialogue between characters. Using the dimensions defined above, we first created a set of all possible combinations of dimensions and then randomly sampled 50 combinations consistent with our crowdsourced weights. Given our sample of dimension combinations, we prompted ChatGPT 4o, Gemini 2.0 Pro, and Claude Sonnet 3.7 to generate the initial iterations of our privacy-relevant scenarios. By generating the scenarios, we hoped to represent a more diverse set of scenario characteristics than we could develop on our own. We next verified that each scenario properly represented the dimensions upon which it was based and made corrections when needed. We additionally edited the scenarios to remove any explicit references to privacy, such that neither the participants nor the studied LLMs would be primed. We present an example script in Figure 1. The full dataset, survey, and the prompt used to generate the scenarios, can be accessed in our OSF repository [^1] https://osf.io/wbp37/?view_only=4a2799ee7cf04480a54d06faf342714d.